OpenStudy (anonymous):

Multiply and simplify. Assume that all variables represent nonnegative numbers. (-5square root 6)( 2square root 3)

OpenStudy (mertsj):

Multiply the coefficients and multiply the radicands

OpenStudy (anonymous):

-10 square root 18

OpenStudy (mertsj):

Now simplify sqrt18

OpenStudy (anonymous):

4.24

OpenStudy (mertsj):

\[-10\sqrt{18}=-10\sqrt{9 \times 2}=-10\sqrt{9}\sqrt{2}=-10 \times 3 \sqrt{2}=-30\sqrt{2}\]